OpenStudy (141823):

What angle is formed by the hands of the clock at 12:05, 7:43, 3:37 MEDALS

OpenStudy (141823):

@Ashleyisakitty Help

OpenStudy (dumbcow):

ok the way i would do this is find the angle the minute hand makes with top of clock or 12:00 then find angle the hour hand makes Finally subtract the two to get angle between the two hands Minute hand: clock is divided into 60 min ---> 360/60 = 6 1 min = 6 degrees whatever number representing minutes, you multiply by 6 to get angle example: 4:30 ----> 30 is the minutes, 30*6 = 180, minute hand has angle of 180 degrees Hour hand: clock is divided into 12 hours --> 360/12 = 30 1 hr = 30 degrees Also we must consider that the hour hand is slowly moving to next hour based on where the minute hand is..... 1 hr is split into 60 min intervals --> 30/60 = 1/2 each minute = 1/2 degree example: 4:30 ---> 4 is hour, 4*30=120 --> 30*1/2 = 15 120 + 15 = 135 hour hand has angle of 135 degrees Using my example of 4:30 min hand = 180 degrees hour hand = 135 subtract ---> 180 - 135 = 45 degrees Also note that if angle is more than 180, subtract from 360 to get correct answer

OpenStudy (141823):

So for 12:05, the minute hand angle is 30 and the hour hand angle is 360???

OpenStudy (141823):

I mean 180

OpenStudy (141823):

But that would mean 180-30=150 and there's no way 12:05 is 150 degrees...

OpenStudy (141823):

Oh wait I forgot it goes from 360 to 180 to 90 and then I subtract 30 from 90 ro get 60 but that still isn't... right.

OpenStudy (dumbcow):

lets look at hour hand again 12:05 12*30 = 360 = 0 (360 means we are back where we started) 5*1/2 = 2.5 0+2.5 = 2.5 hour hand has angle of 2.5 degrees

OpenStudy (141823):

oh, well then. Wait but why is it addition now?

OpenStudy (dumbcow):

??

OpenStudy (141823):

But the angle formed by the hands of the clock at 12:05 can't be 2.5 degrees, the minute and the hour hand form at least a 30 degree angle right?

OpenStudy (141823):

Before you said subtract the minute from the hour or something like that

OpenStudy (dumbcow):

oh no thats just for the hour hand minute hand is still 30 degrees to get angle between them, you still subtract ---> 30 - 2.5 = 27.5

OpenStudy (141823):

But if the hour hand was 360=0 how is it 30?

OpenStudy (141823):

er minute

OpenStudy (dumbcow):

ok at 12:00 both hands are straight up ... angle is 0 as time ticks the minute hand moves faster than hour hand minute hand moves at 6 degrees per minute hour hand only moves at 1/2 degree per minute or 30 degrees per hour

OpenStudy (141823):

Literally can't click Post because of notifications at the bottom -_-

OpenStudy (141823):

But, the 5 minutes in 12:05 accounts for the 2.5 in terms of hour and the 30 in terms of minutes...

OpenStudy (dumbcow):

yes

OpenStudy (141823):

Isn't that like saying x = y but also z but z and y aren't the same??

OpenStudy (dumbcow):

hmm no because there are 2 different hands moving at different speeds

OpenStudy (141823):

Ok I think I get it

OpenStudy (141823):

Let me try another example

OpenStudy (141823):

Clock forms 4:55

OpenStudy (141823):

err Clock is 4:55

OpenStudy (dumbcow):

if the concept is a bit confusing to you i can sum it up in a nice formula given --> H:m \[angl e = |(30H+\frac{1}{2}m) - 6m|\]

OpenStudy (141823):

So 4x30 = 120/2 = 60

OpenStudy (141823):

55/60 = 11/12

OpenStudy (141823):

11/12x30=...

OpenStudy (141823):

dammit

OpenStudy (141823):

Using your formula [(30(4) + 1/2(55) - 6m]

OpenStudy (141823):

or 6(55)?

OpenStudy (dumbcow):

yes, what will that give you?

OpenStudy (141823):

120 + 27.5 - 330 = 147.5 -330

OpenStudy (141823):

I don't even understand

OpenStudy (dumbcow):

dont stop...what is |147.5 -330|?

OpenStudy (141823):

-182.5

OpenStudy (dumbcow):

good, notice i used absolute value to force it to be positive 182.5 the 2 hands cant be more than 180 so in this case we will subtract from 360 360 - 182.5 = 177.5

OpenStudy (141823):

So after the formula if it is - use absolute value and if it's over 180 subtract from 360?

OpenStudy (dumbcow):

correct

OpenStudy (141823):

Okay thanks, let me try one more. See if I can do it on my own. Mind giving me a random time?

OpenStudy (dumbcow):

lets do the ones in your original post 7:43

OpenStudy (141823):

My gosh I can't scroll down to press Post because of the notifications...

OpenStudy (141823):

(30(7) + 1/2(43) - 6(43) 210 + 21.5 - 258 -26.5

OpenStudy (141823):

26.5 is the answer?

OpenStudy (dumbcow):

very good now do 3:37

OpenStudy (dumbcow):

oh yeah its been doing that to me too....press "tab" then "enter" to post without clicking

OpenStudy (141823):

(30(3) + 1/2(37)) - 6(37) 90 + 18.5 -222 108.5-222 -113.5 = 113.5

OpenStudy (dumbcow):

yes looks like you got it, just use the formula i gave you if you want to understand it better you can reread my first post i guess, it may sink in after awhile

OpenStudy (141823):

Yea, it's definitely sinking in

OpenStudy (141823):

Although, I do have one more question but I don't know if I should start a new post for it

OpenStudy (141823):

It has to do with factoring

OpenStudy (dumbcow):

go ahead and just ask me here...i can do a couple more before i log off

OpenStudy (141823):

I know x^2+x-30 is (x-5)(x+6) or something like that but I don't know how to figure it out quickly, like I had a quiz today and the question was x^2 + 11x + 172 = 0 and I blanked :p

OpenStudy (dumbcow):

ok i can give some tips but sometimes, especially with big numbers it can be hard to find the right factors quickly so want 2 numbers that add to middle coefficient and multiply to get last coefficient since 172 is big and have a lot of factors, i would start with the 11 1+10 2+9 3+8 4+7 5+6 and see if any of these multiply to get 172 ... they dont so this is actually not factorable

OpenStudy (141823):

So what would an example with a big number be that would be factorable?

OpenStudy (dumbcow):

x^2 + 32x +240 = 0

OpenStudy (141823):

So I start with 32? Like 31x1 30x2 29x3 etc

OpenStudy (dumbcow):

yeah but some are obviously not going to work like 31 and 1 start in middle 16 +16 17+15 18+14 ...

OpenStudy (141823):

wait 16+16 not 16x16?

OpenStudy (141823):

Jeez I've literally forgotten everything about factoring haha

OpenStudy (dumbcow):

both ... you want it to add to 32 but multiply to 240

OpenStudy (141823):

ok

OpenStudy (141823):

so 16+16 = 32 and 16x16 = 256

OpenStudy (141823):

17+15 = 32 but 17x15 = 255

OpenStudy (141823):

20+12 = 32 but 20x12 = 240

OpenStudy (141823):

ok not what do I do

OpenStudy (141823):

now*

OpenStudy (dumbcow):

yep so its (x+12)(x+20)

OpenStudy (141823):

My God I feel stupid hahahahahaha

OpenStudy (141823):

Silly large numbers

OpenStudy (141823):

What about another example like x^2 +x -30 = 0 where it's (x-5)(x+6) dealing with negatives

OpenStudy (141823):

Are those only for 1x or do I use them for other _x's

OpenStudy (dumbcow):

oh no it could be anything x^2 -5x-14

OpenStudy (141823):

Using my battlefield case to kill mosquitos haha

OpenStudy (141823):

x^2 -5x -14 = 0... (x - 7)(x + 2)?

OpenStudy (dumbcow):

correct

OpenStudy (141823):

yay, I would link an achievement song but I don't know the thread rules on this

OpenStudy (dumbcow):

most links are ok as far as i know....no spam or porn

OpenStudy (141823):

https://www.youtube.com/watch?v=ZOJCmPKaYN8 Murica haha

OpenStudy (141823):

Learned more in one openstudy session than 1 week in Geo, seems legit

OpenStudy (141823):

Alright well I'm definitely done for the night. Thanks for all your help!

OpenStudy (dumbcow):

hey yw feel free to tag me if im on if you need more help

OpenStudy (141823):

ok cool